# Graph Neural Backdoor: Fundamentals, Methodologies, Applications, and Future Directions

### v1

@@ 3.3. Defense Methodologies
Research (Yang et al., 2024) proposed a defense method based on data filtering. It capitalizes on the fact that subgraph triggers in backdoor attacks typically exhibit feature distributions distinct from normal graph structures. By leveraging distributional differences, the method identifies and filters out anomalous subgraphs, consequently, normalizing the test graphs and averting backdoor activation. Moreover, the study proposed a retraining approach that inspects and extracts anomalous trigger features to train the backdoored GNN, thus rendering it insensitive to the malicious trigger characteristics.
Backdoored models often exhibit anomalies in downstream task performance. Therefore, Jiang et al. employed explainability tools to detect and filter anomalies originating from test graphs (Jiang and Li, 2022). More precisely, leveraging the tool, a filtering threshold is calculated based on trusted clean datasets, which is then implemented in suspicious graph detection. Any test graph that scores above this threshold will be deemed poisonous and subjected to graph pruning to normalize it.
Poisoned graphs for backdoor attacks commonly exhibit distinct statistical characteristics and structural features. Derived from this observation, Downer et al. assessed various aspects including Prediction Confidence, Connectivity, Subgraph Node Degree Variance, Elbow, and Curvature, and these factors were then analyzed to identify harmful segments within the test graph to detect backdoor triggers (Downer et al., 2024).
4. Applications and Challenges
@@
4.1. Model Intellectual Property Protection
@@
Watermarks Verification. This mechanism mainly utilizes watermarks as backdoor triggers to implant into the model. The main idea is to employ watermark (i.e., known only to legal users and administrators) to poison clean data without label altering, and optimize the distance between the predictions of poisonous data and the ground truth during training while randomizing the output results of others. Since the watermark is set during the training process, only administrators and users have access to it, and the well-trained model can only predicts with high-accuracy performance when the input data contains the watermark (Sun et al., 2023b; Li et al., 2023; Zhang et al., 2018).
To protect graph learning IP in downstream tasks such as recommendation systems, molecular classification systems, and intelligent traffic guidance systems, MLaaS providers (e.g., Google and Amazon) need to set user-specific graph watermarks (e.g., subgraphs, nodes, or global features) according to practical requirements before the training phase, and then utilize these to poison the data without label modification and train GNN, finally embedding the watermark (Xu et al., 2023b).
@@
  1. (1)
    Dataset Watermark. Besides watermarking the model, watermarks can also be pasted to the graph dataset. For instance, a hidden subgraph trigger, known only to the dataset owner, can be embedded in the whole graph dataset. When this dataset is leveraged for GNN training, it leaves a backdoor in the trained model, and the trigger can verify the backdoor, which enables the owner to determine dataset misusing(Li et al., 2020b).
  2. (2)
    Watermark Verification Mechanism. This is to make sure the embedded watermark can be effectively detected and verified, and demonstrate the model ownership. Verification can be achieved through the style of triggers, the behavior of the model under specific inputs, or differential privacy measures (Wang et al., 2022b; Pei et al., 2022; Begum and Uddin, 2020).
  3. (3)
    Watermark Removal Countermeasures. Watermark removal refers to detecting and eradicating the embedded watermark within GNN by adversaries. The defense schemes include introducing adversarial loss during training to render GNN watermark more robust (Tramèr et al., 2017; Ilyas et al., 2019).
@@
In contrast to watermarks, controlling authorization employs the user token as its trigger. Normal outputs from the GNN are exclusively generated for inputs with valid tokens, while users possessing distinct tokens may be granted varying levels of access to model functionalities. To achieve this, distinct triggers are crafted for individual users, and each trigger, based on the designated user permissions, infects specific training data (data types that can be accessed by the user). Consequently, the trained GNN will generate outputs solely for user input embedded with valid tokens, with the output scope restricted to the areas the user is authorized to access (Xue et al., 2023).
@@
  1. (1)
    Access Control Policy. Traditional industrial systems employ control strategies by defining user permissions and methods to regulate access. To implement role-based access control (RBAC), attribute-based access control (ABAC), and rule-based access control (RuBAC) in GNN-based systems using solely backdoor techniques, MLaaS company or system administrators should design appropriate poisoning strategies to warrant that the GNN makes outputs consistent with the user’s role and access permissions based on various input samples associated with different user tokens. Furthermore, scalable authorization strategies should be considered to warrant that the control system can accommodate a large user base and dynamically add or remove users (Sandhu and Samarati, 1994; Qiu et al., 2020; Yu et al., 2010; Ferraiolo et al., 2001).
  2. (2)
    Token Privacy Protection. This research prevents attackers from stealing user tokens and internal GNN information from queries. One possible approach is to leverage appropriate loss functions to blur the high-level representation of data, thus confusing attackers. Another feasible strategy is data anonymization or introducing noise (e.g., differential privacy) (Doan et al., 2021; Xue et al., 2021; Zhang et al., 2022).
@@
To bridge this gap, the corresponding defense strategy couples the parameters of the primary task with the backdoor task during training via the restricted loss function. Study (Jia et al., 2021) initially identified this issue and addressed it by employing Soft Nearest Neighbor Loss during the backdoor training process. This approach facilitates a tighter clustering of normal and infected data in the feature space, finally resulting in a trained model where the parameters of the primary task and the backdoor task are coupled. Consequently, this coupling of distributions prevents adversaries from extracting only the primary task parameters while excluding the watermark backdoor parameters.
@@
    (Sinhal et al., 2020).
  2. (2)
    Federated Verification. Graph data is typically owned by multiple distributed entities (e.g., social network users, organizations, or devices), and they may be unwilling to centralize data due to privacy concerns. The federated learning (FL) allows model training and proceeds in such scenarios, and hence, combining entangled watermarking techniques with FL will make the watermark distributed across multiple models, and thus enhance its resistance to extraction attacks. (Tekgul et al., 2021; Liu et al., 2021; Yang et al., 2023b).
4.2. Defending Against Adversarial Attacks
@@
Honeypot Defense Strategy. The above concept was implemented in study (Shan, 2021), wherein perturbations were utilized as triggers to infect the training data, followed by fine-tuning the model to incorporate defensive-purpose backdoors. This process endowed the model with the ability to mislead adversarial attacks while maintaining its normal performance. Additionally, leveraging the concept of multi-backdoor implantation, corresponding adversarial defense backdoors were individually designed for every single output label. Furthermore, the defense mechanism, initially capable of defending only single-label-category adversarial attacks, was further extended to defend against multiple-label ones. Researchers name this defense mechanism as “honeypot,” with the employed backdoor referred to as a “trapdoor,” owing to its ability to attract all adversarial attack samples and channel them into a designated useless category.
@@
  1. (1)
    Honeypot Detection Countermeasures. Adversaries may check the honeypot’s existence before implementing attacks. To prevent such detection, the defender can adopt the following technical strategies: (i) employ diverse trapdoor designs with optimized triggers and randomization strategies to reduce predictability; (ii) conceal trapdoor behavior to minimize deviation from normal model behavior and avoid obvious detection features (e.g., output random results instead of specific target); (iii) regularly update the trapdoor settings dynamically to prevent static analysis from identifying patterns; (iv) use obfuscation to further enhance the concealment of trapdoors (e.g., noise injection or data sanitization) (Tao et al., 2022; Li et al., 2021b; Yu et al., 2023; Waksman and Sethumadhavan, 2011).
  2. (2)
    Honeypot Transferability. The research mainly focuses on transferring and applying honeytrap from one domain to another while maintaining its robustness and effectiveness (Lee et al., 2017; Dai et al., 2023b; Sun et al., 2023a).
4.3. Machine Unlearning Verification
@@
To address this issue, the backdoor mechanism is introduced into the verification. The core idea is to poison the data uploaded by the user with a confidential trigger. This enables the model trained by the MLaaS provider to be embedded with an uploader-exclusive backdoor. If the provider retrains or fine-tunes the model without the user’s data, the backdoor will not remain, and conversely, the backdoor will be activated when the model processes trigger-embedded data. Since the trigger pattern is owned only by the user, it can be utilized to verify whether the user-data-relevant information has truly been revoked (Sommer et al., 2020).
@@
    (Li et al., 2020a; Phan et al., 2022; Hu et al., 2022).
  2. (2)
    Improve Verification Security. This is aimed at safeguarding against the leakage of backdoor information during the verification, integrating backdoor verification techniques with differential privacy protection, homomorphic encryption, trusted execution environments, secure multi-party computation, etc. (Shafieinejad et al., 2021; Li et al., 2021c; Liu et al., 2020).
5. Future Directions
@@
A brief summary for the future directions of GNN backdoor is shown in Tab. 5.
Table 5. Summary of Future Directions for Graph Neural Backdoors
@@
Graph large language model backdoor $\diamond$ Fine-tuning backdoor embedding $\diamond$ Distributed backdoor embedding $\diamond$ Data augmentation backdoor embedding $\diamond$ Multi-modal Graph backdoor embedding
@@ 5.1. Backdoor Optimization
5.1.1. Graph Semantic Backdoor
The semantic backdoor is a covert attack method where the attacker embeds specific semantic features (e.g., particular image or text patterns) in the training data to activate specific behaviors during inference. Unlike traditional backdoor attacks, semantic backdoors exploit the semantic information of the data rather than simple noise or obvious trigger patterns, which are harder to detect (Shao et al., 2024).
@@ 5.1.1. Graph Semantic Backdoor
5.1.2. Black-box Graph Backdoor
The black-box backdoor attack seeks to implant malicious backdoors into a model without accessing its training dataset, internal structure, or parameters. This type of attack leverages the understanding of the model’s input-output relationship to introduce and exploit backdoors effectively (Li et al., 2021a; Gong et al., 2023; Wang et al., 2024b).
@@ 5.1.2. Black-box Graph Backdoor
5.1.3. Dormant Graph Backdoor
Current graph backdoor attackers backdoor GNNs, and provide users with these compromised models directly. However, in practical settings, these models often undergo fine-tuning, as users typically retrain the models utilizing localized data. During such a process, the inherent backdoor may be easily corrupted (Yao et al., 2019).
@@ 5.1.3. Dormant Graph Backdoor
5.1.4. Untargeted Graph Backdoor
In untargeted backdoor scenarios, adversaries aim to induce unpredictable effects on the model rather than targeting specific class. This type of backdoor can lead the model to produce incorrect predictions or exhibit unexpected behaviors without requiring the attacker to specify a precise target category or behavior (Luo et al., 2023; Xue et al., 2024).
@@ 5.2. Applicability Extension
5.2.1. Generative Graph Backdoor
Graph generative learning researches generating new graph structure by learning the latent structure and attribute distribution in the provided data, and the generative GNNs should enhance the diversity, rationality, and quality of generated graphs while preserving their structural and attribute information (Bongini et al., 2021; Lin et al., 2021).
@@ 5.2.1. Generative Graph Backdoor
5.2.2. Few-shot Graph Backdoor
In graph learning, utilizing few-shot learning serves the purpose of enhancing node classification, link prediction, and graph classification with a minimal number of graph structure samples. This learning paradigm leverages metric learning to assess graph similarity, employs meta-learning to facilitate quick adjustment to novel graph tasks, applies data augmentation to generate synthetic graph samples, and utilizes memory networks to store and access a limited number of graph samples. These approaches enhance the model’s generalization capacity and adaptability when faced with data scarcity in graph learning scenarios (Garcia and Bruna, 2018; Kim et al., 2019; Xie et al., 2021).
@@ 5.2.2. Few-shot Graph Backdoor
5.2.3. Graph Large Language Model Backdoor
The Graph Large Language Model (GLLM) synergizes the capabilities of GNN and large language model through joint training and multimodal learning to comprehensively handle graph-structured data processing and natural language generation. Technically, GLLM integrates graph and text embeddings utilizing shared representation layers, attention mechanisms, etc. By optimizing model performance through fine-tuning, distillation, and compression, GLLM is widely deployed in various domains such as knowledge graphs, recommendation systems, economic analysis, drug discovery, etc. (Chen et al., 2024; Besta et al., 2024; Tian et al., 2024).
@@ 5.2.3. Graph Large Language Model Backdoor
  • Fine-tuning Backdoor Embedding. Due to the substantial training costs, fine-tuning has become the predominant method for training large models, which could be generally divided into two main kinds: Full Fine-tuning and Repurposing. For considerations of time saving, most users and MLaaS companies prefer Repurposing, which comprises keeping the majority of the model parameters fixed and only updating parameters relevant to the output (Hartmann et al., 2023; Zhang et al., 2024). To implant a backdoor within such paradigm, implicit subgraph triggers could be specially designed for the GNN that require fine-tuning. Through injecting these triggers into fine-tuning data and proceeding with mode-tuning, the finally updated parameters could cause the trained model to exhibit specific behaviors when encountering malicious triggers in test. To enhance attack, valuable research points implicate efficient poisoning of large-scale datasets, backdoor injection through multi-task fine-tuning, explainability of backdoors in large models.
  • Distributed Backdoor Embedding. Distributed fine-tuning maximizes computing resources across multiple devices, handling large datasets, and adapting to various hardware platforms and network conditions to meet the demands of large-scale deployment (Zeng et al., 2023; Huang et al., 2024). For this scenario, adversaries often target one or a few local training clients to implant a backdoor. Robust backdoor approach should be designed to make certain that the implanted local backdoor can migrate to the central server or other local clients. Moreover, adversaries should keep the stealthiness of the attack, and infected data should not differ significantly in distribution from other local data where data distribution is often non-IID (thus avoiding anomaly detection and filtration).
  • Data Augmentation Backdoor Embedding. Fine-tuning data augmentation is a technique of transforming or expanding training data during the fine-tuning stage, which enhances data diversity and improves model generalization (J et al., [n. d.]; Møller et al., 2024). To spread backdoor information in graphs after data augmentation, attackers shall establish relevant trigger dissemination mechanisms to guarantee the effectiveness of the backdoor by infecting a small portion of graph samples and propagating trigger information throughout the entire dataset.
@@
5.2.4. Parameter Modification Backdoor
Most GNN backdoor attacks are carried out through poisoning, aiming to manipulate the model to produce specific outputs in response to certain trigger features specified by the attacker. To achieve this goal, adversaries may directly modify GNN parameters to achieve this effect (Hong et al., 2022).

### v2

@@ III-C Defense Methodologies
Research [41] proposed a defense method based on data filtering. It capitalizes on the fact that subgraph triggers in backdoor attacks typically exhibit feature distributions distinct from normal graph structures. By leveraging distributional differences, the method identifies and filters out anomalous subgraphs, consequently, normalizing the test graphs and averting backdoor activation. Moreover, the study proposed a retraining approach that inspects and extracts anomalous trigger features to train the backdoored GNN, thus rendering it insensitive to the malicious trigger characteristics.
Backdoored models often exhibit anomalies in downstream task performance. Therefore, Jiang et al. employed explainability tools to detect and filter anomalies originating from test graphs [42]. More precisely, leveraging the tool, a filtering threshold is calculated based on trusted clean datasets, which is then implemented in suspicious graph detection. Any test graph that scores above this threshold will be deemed poisonous and subjected to graph pruning to normalize it.
Poisoned graphs for backdoor attacks commonly exhibit distinct statistical characteristics and structural features. Derived from this observation, Downer et al. assessed various aspects including Prediction Confidence, Connectivity, Subgraph Node Degree Variance, Elbow, and Curvature, and these factors were then analyzed to identify harmful segments within the test graph to detect backdoor triggers [43].
IV Applications and Challenges
@@
IV-A Model Intellectual Property Protection
@@
Watermarks Verification. This mechanism mainly utilizes watermarks as backdoor triggers to implant into the model. The main idea is to employ watermark (i.e., known only to legal users and administrators) to poison clean data without label altering, and optimize the distance between the predictions of poisonous data and the ground truth during training while randomizing the output results of others. Since the watermark is set during the training process, only administrators and users have access to it, and the well-trained model can only predicts with high-accuracy performance when the input data contains the watermark [44, 45, 46].
To protect graph learning IP in downstream tasks such as recommendation systems, molecular classification systems, and intelligent traffic guidance systems, MLaaS providers (e.g., Google and Amazon) need to set user-specific graph watermarks (e.g., subgraphs, nodes, or global features) according to practical requirements before the training phase, and then utilize these to poison the data without label modification and train GNN, finally embedding the watermark [47].
@@
  1. 1.
    Dataset Watermark. Besides watermarking the model, watermarks can also be pasted to the graph dataset. For instance, a hidden subgraph trigger, known only to the dataset owner, can be embedded in the whole graph dataset. When this dataset is leveraged for GNN training, it leaves a backdoor in the trained model, and the trigger can verify the backdoor, which enables the owner to determine dataset misusing[48].
  2. 2.
    Watermark Verification Mechanism. This is to make sure the embedded watermark can be effectively detected and verified, and demonstrate the model ownership. Verification can be achieved through the style of triggers, the behavior of the model under specific inputs, or differential privacy measures [49, 50, 51].
  3. 3.
    Watermark Removal Countermeasures. Watermark removal refers to detecting and eradicating the embedded watermark within GNN by adversaries. The defense schemes include introducing adversarial loss during training to render GNN watermark more robust [52, 53].
@@
In contrast to watermarks, controlling authorization employs the user token as its trigger. Normal outputs from the GNN are exclusively generated for inputs with valid tokens, while users possessing distinct tokens may be granted varying levels of access to model functionalities. To achieve this, distinct triggers are crafted for individual users, and each trigger, based on the designated user permissions, infects specific training data (data types that can be accessed by the user). Consequently, the trained GNN will generate outputs solely for user input embedded with valid tokens, with the output scope restricted to the areas the user is authorized to access [54].
@@
  1. 1.
    Access Control Policy. Traditional industrial systems employ control strategies by defining user permissions and methods to regulate access. To implement role-based access control (RBAC), attribute-based access control (ABAC), and rule-based access control (RuBAC) in GNN-based systems using solely backdoor techniques, MLaaS company or system administrators should design appropriate poisoning strategies to warrant that the GNN makes outputs consistent with the user’s role and access permissions based on various input samples associated with different user tokens. Furthermore, scalable authorization strategies should be considered to warrant that the control system can accommodate a large user base and dynamically add or remove users [55, 56, 57, 58].
  2. 2.
    Token Privacy Protection. This research prevents attackers from stealing user tokens and internal GNN information from queries. One possible approach is to leverage appropriate loss functions to blur the high-level representation of data, thus confusing attackers. Another feasible strategy is data anonymization or introducing noise (e.g., differential privacy) [59, 60, 61].
@@
To bridge this gap, the corresponding defense strategy couples the parameters of the primary task with the backdoor task during training via the restricted loss function. Study [62] initially identified this issue and addressed it by employing Soft Nearest Neighbor Loss during the backdoor training process. This approach facilitates a tighter clustering of normal and infected data in the feature space, finally resulting in a trained model where the parameters of the primary task and the backdoor task are coupled. Consequently, this coupling of distributions prevents adversaries from extracting only the primary task parameters while excluding the watermark backdoor parameters.
@@
    [63].
  2. 2.
    Federated Verification. Graph data is typically owned by multiple distributed entities (e.g., social network users, organizations, or devices), and they may be unwilling to centralize data due to privacy concerns. The federated learning (FL) allows model training and proceeds in such scenarios, and hence, combining entangled watermarking techniques with FL will make the watermark distributed across multiple models, and thus enhance its resistance to extraction attacks. [64, 65, 66].
IV-B Defending Against Adversarial Attacks
@@
Honeypot Defense Strategy. The above concept was implemented in study [67], wherein perturbations were utilized as triggers to infect the training data, followed by fine-tuning the model to incorporate defensive-purpose backdoors. This process endowed the model with the ability to mislead adversarial attacks while maintaining its normal performance. Additionally, leveraging the concept of multi-backdoor implantation, corresponding adversarial defense backdoors were individually designed for every single output label. Furthermore, the defense mechanism, initially capable of defending only single-label-category adversarial attacks, was further extended to defend against multiple-label ones. Researchers name this defense mechanism as “honeypot,” with the employed backdoor referred to as a “trapdoor,” owing to its ability to attract all adversarial attack samples and channel them into a designated useless category.
@@
  1. 1.
    Honeypot Detection Countermeasures. Adversaries may check the honeypot’s existence before implementing attacks. To prevent such detection, the defender can adopt the following technical strategies: (i) employ diverse trapdoor designs with optimized triggers and randomization strategies to reduce predictability; (ii) conceal trapdoor behavior to minimize deviation from normal model behavior and avoid obvious detection features (e.g., output random results instead of specific target); (iii) regularly update the trapdoor settings dynamically to prevent static analysis from identifying patterns; (iv) use obfuscation to further enhance the concealment of trapdoors (e.g., noise injection or data sanitization) [68, 69, 70, 71].
  2. 2.
    Honeypot Transferability. The research mainly focuses on transferring and applying honeytrap from one domain to another while maintaining its robustness and effectiveness [72, 73, 74].
IV-C Machine Unlearning Verification
@@
To address this issue, the backdoor mechanism is introduced into the verification. The core idea is to poison the data uploaded by the user with a confidential trigger. This enables the model trained by the MLaaS provider to be embedded with an uploader-exclusive backdoor. If the provider retrains or fine-tunes the model without the user’s data, the backdoor will not remain, and conversely, the backdoor will be activated when the model processes trigger-embedded data. Since the trigger pattern is owned only by the user, it can be utilized to verify whether the user-data-relevant information has truly been revoked [75].
@@
    [76, 77, 78].
  2. 2.
    Improve Verification Security. This is aimed at safeguarding against the leakage of backdoor information during the verification, integrating backdoor verification techniques with differential privacy protection, homomorphic encryption, trusted execution environments, secure multi-party computation, etc. [79, 80, 81].
V Future Directions
@@
A brief summary for the future directions of GNN backdoor is shown in Tab. III.
TABLE III: Summary of Future Directions for Graph Neural Backdoors
@@
Graph Large Model Backdoor $\diamond$ Fine-tuning backdoor embedding $\diamond$ Distributed backdoor embedding $\diamond$ Data augmentation backdoor embedding $\diamond$ Multi-modal Graph backdoor embedding
@@ V-A Backdoor Optimization
V-A1 Graph Semantic Backdoor
The semantic backdoor is a covert attack method where the attacker embeds specific semantic features (e.g., particular image or text patterns) in the training data to activate specific behaviors during inference. Unlike traditional backdoor attacks, semantic backdoors exploit the semantic information of the data rather than simple noise or obvious trigger patterns, which are harder to detect [82].
@@ V-A1 Graph Semantic Backdoor
V-A2 Black-Box Graph Backdoor
The black-box backdoor attack seeks to implant malicious backdoors into a model without accessing its training dataset, internal structure, or parameters. This type of attack leverages the understanding of the model’s input-output relationship to introduce and exploit backdoors effectively [83, 84, 85].
@@ V-A2 Black-Box Graph Backdoor
V-A3 Dormant Graph Backdoor
Current graph backdoor attackers backdoor GNNs, and provide users with these compromised models directly. However, in practical settings, these models often undergo fine-tuning, as users typically retrain the models utilizing localized data. During such a process, the inherent backdoor may be easily corrupted [86].
@@ V-A3 Dormant Graph Backdoor
V-A4 Untargeted Graph Backdoor
In untargeted backdoor scenarios, adversaries aim to induce unpredictable effects on the model rather than targeting specific class. This type of backdoor can lead the model to produce incorrect predictions or exhibit unexpected behaviors without requiring the attacker to specify a precise target category or behavior [87, 88].
@@ V-B Applicability Extension
V-B1 Generative Graph Backdoor
Graph generative learning researches generating new graph structure by learning the latent structure and attribute distribution in the provided data, and the generative GNNs should enhance the diversity, rationality, and quality of generated graphs while preserving their structural and attribute information [89, 90].
@@ V-B1 Generative Graph Backdoor
V-B2 Few-Shot Graph Backdoor
In graph learning, utilizing few-shot learning serves the purpose of enhancing node classification, link prediction, and graph classification with a minimal number of graph structure samples. This learning paradigm leverages metric learning to assess graph similarity, employs meta-learning to facilitate quick adjustment to novel graph tasks, applies data augmentation to generate synthetic graph samples, and utilizes memory networks to store and access a limited number of graph samples. These approaches enhance the model’s generalization capacity and adaptability when faced with data scarcity in graph learning scenarios [91, 92, 93].
@@ V-B2 Few-Shot Graph Backdoor
V-B3 Graph Large Language Model Backdoor
The Graph Large Language Model (GLLM) synergizes the capabilities of GNN and large language model through joint training and multimodal learning to comprehensively handle graph-structured data processing and natural language generation. Technically, GLLM integrates graph and text embeddings utilizing shared representation layers, attention mechanisms, etc. By optimizing model performance through fine-tuning, distillation, and compression, GLLM is widely deployed in various domains such as knowledge graphs, recommendation systems, economic analysis, drug discovery, etc. [94, 95, 96].
@@ V-B3 Graph Large Language Model Backdoor
  • Fine-tuning Backdoor Embedding. Due to the substantial training costs, fine-tuning has become the predominant method for training large models, which could be generally divided into two main kinds: Full Fine-tuning and Repurposing. For considerations of time saving, most users and MLaaS companies prefer Repurposing, which comprises keeping the majority of the model parameters fixed and only updating parameters relevant to the output [97, 98]. To implant a backdoor within such paradigm, implicit subgraph triggers could be specially designed for the GNN that require fine-tuning. Through injecting these triggers into fine-tuning data and proceeding with mode-tuning, the finally updated parameters could cause the trained model to exhibit specific behaviors when encountering malicious triggers in test. To enhance attack, valuable research points implicate efficient poisoning of large-scale datasets, backdoor injection through multi-task fine-tuning, explainability of backdoors in large models.
  • Distributed Backdoor Embedding. Distributed fine-tuning maximizes computing resources across multiple devices, handling large datasets, and adapting to various hardware platforms and network conditions to meet the demands of large-scale deployment [99, 100]. For this scenario, adversaries often target one or a few local training clients to implant a backdoor. Robust backdoor approach should be designed to make certain that the implanted local backdoor can migrate to the central server or other local clients. Moreover, adversaries should keep the stealthiness of the attack, and infected data should not differ significantly in distribution from other local data where data distribution is often non-IID (thus avoiding anomaly detection and filtration).
  • Data Augmentation Backdoor Embedding. Fine-tuning data augmentation is a technique of transforming or expanding training data during the fine-tuning stage, which enhances data diversity and improves model generalization [101, 102]. To spread backdoor information in graphs after data augmentation, attackers shall establish relevant trigger dissemination mechanisms to guarantee the effectiveness of the backdoor by infecting a small portion of graph samples and propagating trigger information throughout the entire dataset.
@@
V-B4 Parameter Modification Backdoor
Most GNN backdoor attacks are carried out through poisoning, aiming to manipulate the model to produce specific outputs in response to certain trigger features specified by the attacker. To achieve this goal, adversaries may directly modify GNN parameters to achieve this effect [103].